OpenStudy (anonymous):

A bag contains six purple marbles and 7 white marbles. two marbles are drawn at random. one marble is drawn but not replaced. then a second marble is drawn. What is the probability that the first marble is white and the second purple?

OpenStudy (anonymous):

=(Prob 1st white)*(prob 2nd purple) =(7/13)*(6/12) =(7/13)*(1/2) =7/26

OpenStudy (anonymous):

Thank You!(: